# FinCon: A Synthesized LLM Multi-Agent System with Conceptual Verbal Reinforcement for Enhanced Financial Decision Making

### v2

@@ Abstract
Large language models (LLMs) have demonstrated notable potential in conducting complex tasks and are increasingly utilized in various financial applications. However, high-quality sequential financial investment decision-making remains challenging. These tasks require multiple interactions with a volatile environment for every decision, demanding sufficient intelligence to maximize returns and manage risks. Although LLMs have been used to develop agent systems that surpass human teams and yield impressive investment returns, opportunities to enhance multi-sourced information synthesis and optimize decision-making outcomes through timely experience refinement remain unexplored. Here, we introduce the FinCon, an LLM-based multi-agent framework with conceptual verbal reinforcement tailored for diverse financial tasks. Inspired by effective real-world investment firm organizational structures, FinCon utilizes a manager-analyst communication hierarchy. This structure allows for synchronized cross-functional agent collaboration towards unified goals through natural language interactions and equips each agent with greater memory capacity than humans. Additionally, a risk-control component in FinCon enhances decision quality by episodically initiating a self-critiquing mechanism to update systematic investment beliefs. The conceptualized beliefs serve as verbal reinforcement for the future agent’s behavior and can be selectively propagated to the appropriate node that requires knowledge updates. This feature significantly improves performance while reducing unnecessary peer-to-peer communication costs. Moreover, FinCon demonstrates strong generalization capabilities in various financial tasks, including single stock trading and portfolio management.
@@ 1 Introduction
The intricacies and fluctuations of financial markets present substantial challenges for making high-quality, sequential investment decisions. In tasks such as single stock trading and portfolio management, each intelligent decision is informed by multiple market interactions and integrated information characterized by different levels of timeliness and modalities [1, 2]. These tasks aim to maximize profit while managing current market risks in an open-end environment.
@@
In practice, trading firms often rely on synthesized human teamwork, with an organizational structure that typically involves hierarchical communication among various functional roles such as data analysts, risk analysts, portfolio managers, etc., and careful integration of various resources [3, 4]. However, the cognitive limitations of human team members can restrict their ability to quickly process market signals and achieve optimal investment results [5].
@@
To improve investment returns and overcome human limitations, studies have employed methods such as deep reinforcement learning (DRL) to develop agent systems that simulate market environments and automate investment strategies [6, 7, 8]. Meanwhile, advancements in large language models (LLMs) demonstrate potential in complex tasks such as reasoning, tool-using, planning, and decision-making [9, 10], promising to surpass existing agent architectures. Language agents are known for human-like communication and adaptable prompt-based structures tailored for diverse decision-making settings [11, 12, 13, 14]. For optimal decision-making performance, two key elements need to be considered: 1) Organizing agents to promote effective teamwork and efficient communication, and 2) Enabling agents to continually learn and refine their actions. Studies indicate that imitating human organizational structures can effectively coordinate language agents for specific tasks [15, 16, 17]. Additionally, recent developments in textual gradient-based prompt optimization [18, 19] and verbal reinforcement [20, 21] have proven to be practical for enhancing the reasoning and decision-making abilities of language agents iteratively.
@@
Language agent systems adapted for financial decision-making, such as FinGPT [22], FinMem [23] and FinAgent [24], have demonstrated impressive performance. However, these systems exhibit several limitations. Firstly, their reliance on the risk preferences of agents based on short-term market fluctuations fails to control overall risk exposures over long-term trading periods. This oversight can lead to potential losses due to the dismissal of fundamental factors driving investment profits. A more effective strategy is to quantify investment risks using measures of risk from quantitative finance [25, 26]. Secondly, most existing systems are designed for single-asset trading tasks and show limited adaptability to other financial applications involving multiple assets, such as portfolio management. Thirdly, these systems depend on the information understanding and extraction ability of a single agent, placing heavy pressure on the language model to comprehend and process information within a limited context window. This likely degrades decision quality. A multi-agent system that allocates tasks based on data source and modality could enhance performance. Even though approaches like StockAgent [27] utilize a multi-agent system for individual stock trading tasks. However, their decision-making processes heavily depend on extensive discussions among numerous LLM agents, leading to high communication costs and extended processing times. Besides, the lack of a clear optimization goal in such systems can compromise the effectiveness of the outcomes.
@@
In response to these unaddressed issues, we propose FinCon, an LLM-based multi-agent framework for critical financial decision-making tasks such as single stock trading and portfolio management. Our main contributions are as follows: 1) Inspired by real-world investment roles, we designed an innovative Synthesized Manager-Analyst hierarchical communication structure and a risk control component. This communication structure distributes financial data from various sources and modalities to corresponding functional analyst agents, allowing them to focus on distilling essential investment insights and indicators from a single source of information. The manager then makes trading decisions by consolidating these insights and interacting with the risk-control component. The manager and analyst in a synchronous hierarchy to effectively enhance overall comprehension and reasoning capabilities. In this communication network, messages are conveyed only to the workers who need to be informed, eliminating redundant peer-to-peer communication and saving communication costs. 2) It is a generalized framework that can conduct not only stock trading but also portfolio management. The latter has not been claimed to be tackled by other financial language agent frameworks. 3) We innovatively crafted the agents’ persona using a dual-level risk control component, enabling comprehensive risk updates both within and across episodes. Within episodes, risk supervision is conducted using a quantile-based risk measure from quantitative finance, the Conditional Value at Risk (CVaR) [28]. For across-episode risk control, we designed a unique verbal reinforcement method to update the decision maker’s investment beliefs with conceptualized investment insights to guide future decisions. These investment beliefs are distilled from underlying reasoning trajectory-level profit and loss (PnL) changes and summarized as conceptual perspectives based on the functionality of all analyst agents. Investment beliefs can then be selectively back-propagated from the manager to the corresponding functional analyst agent. Ablation study results illustrate the significant effectiveness of our risk control design in managing market risk and improving trading performance.
2 Related Work
LLM Agents for Financial Decision Making. There are considerable efforts directed towards developing general-purpose LLM agent for sequential decision-making [29, 30], and such type of tasks often involve episodic interactions with environment and verbal reflections for action refinement, such as coding competition [31, 32], software development [33, 14], game-playing [34, 35]. Furthermore, researchers have started to exploit how LLM agents can perform better in harder decision-making tasks from finance [36, 37, 38], in which there are more volatile environments, leading to that the numerous unpredictable elements can obscure an agent’s ability to reflect accurately on the reasons for poor decision outcomes. FinMem [23] enhances single stock trading performance by embedding memory modules with LLM agent for reflection-refinement, and FinAgent [24] improved trading profits via using external quantitative tool to fight against volatile environment.
Multi-Agent System and Communication Structures. In traditional multi-agent systems [39, 40], the way for agents’ communication is pre-determined, like sharing data or state observations [41, 42, 43, 44, 45]. The emergence of large language model brings flexibility for human-understandable communications [46, 11, 14, 47], so some work tries to elevate decision-making ability of LLM-based multi-agent system by letting agents engage in discussions [48, 12] or debates [49, 50]. The similar peer-communication strategy was as well utilized by the multi-agent system for financial tasks [51, 52, 53]. However, such approach are not optimal for unified-goal financial tasks that prioritize profits [54], because they suffer from potentially ambiguous optimization objectives and are unable to control the unnecessary communication costs [55].
Prompt Optimization and Verbal Reinforcement. To enhance the reasoning or decision-making of LLM agents, many prompt optimization techniques have been proposed, like ReAct [56], Chain of Thought (CoT) [57], Tree of Thoughts (ToT) [58], ART [10], intended for that LLM agents can automatically generate intermediate reasoning steps as an iterative program. In addition, to make LLM agent make decisions like humans and generate more understandable reasoning texts, some researchers recommend incorporating cognitive structures [59, 60, 61, 62]. Inspired by these previous work and DRL algorithms [63], verbal reinforcement [20, 21, 64, 15] was developed for LLM agents such that they can update actions based on iterative self-reflection while integrating additional LLM as a prompt optimizer [18, 19].
@@ 3.1 Financial Decision-making Tasks Formulation
Single Stock Trading Tasks. FinCon uses analyst agents group $\{M_{pr}^{i}\}_{i=1}^{I}$ to process multi-modal market information sources. The processed information is then used by a manager agent $M_{a}$ to make trading decisions (long, short, neutral), and to provide relevant reasoning texts. Note that the “short” signal means we make a “short-selling” decision, that is, a negative trading position is allowed. Additionally, FinCon evaluates the daily investment risk, followed by prompt-optimization for the manager agent from risk-control component $M_{r}$.
Portfolio Trading Tasks. In addition to processing multi-modal market information, the analyst agents also establish stocks’ pool for portfolio management, regarding statistical correlation between stocks’ returns. Then, the manager agent makes trading decisions on each stock in the pool. Lastly, the manager agent determines the portfolio weights for all stocks through a computational tool, which leverages the following mean-variance optimization [66]:
$\max_{\textbf{w}}\langle\textbf{w},\mu\rangle-\langle\textbf{w},\Sigma\textbf{w}\rangle\quad\text{s.t.}~{}w_{n}=\begin{cases}\in[0,1],&\text{``long''}\\ \in[-1,0],&\text{``short''}\\ =0,&\text{``neutral''}\end{cases},~{}~{}\forall n\in\{1,\cdots,N\}$
where $\textbf{w}=(w_{1}\cdots,w_{N})\in\mathbb{R}^{N}$ is portfolio weights vector, $\mu$ and $\Sigma$ are the shrinkage estimators of $N$-dimensional sample expected return and $N\times N$ sample covariance matrix of chosen stocks’ daily return sequences respectively [26]. We note that portfolio weights are rebalanced on daily basis. In our implementation, we first calculate the portfolio weights by solving the above optimization problem. Then, the target position is calculated by linearly scaling the portfolio weights from the prior step.
@@ 3.2 Modeling Quantitative Trading as POMDP
Then, our multi-agent system is supposed to learn the policies of all agents: the policies of analyst agents $\pi_{\theta^{i}}^{i}:\mathcal{X}\to\mathcal{A}^{i},i\in\{1,\cdots,I\}$ (the ways to process information, i.e. $\hat{O}_{t}^{i}\sim\pi_{\theta^{i}}^{i}(\cdot|O_{t}^{i})$), and the policy of manager agent $\pi_{\theta^{a}}:\mathcal{A}\times\mathcal{Y}\to\mathbb{A}$ (the ways to make trading decisions, i.e. $A_{t}\sim\pi_{\theta^{a}}(\cdot|\hat{O}_{t},B_{t})$) such that the system maximizes cumulative trading reward while controlling risk [70].
All policies $\Pi_{\bm{\theta}}=(\{\pi_{\theta^{i}}^{i}\}_{i=1}^{I},\pi_{\theta^{a}})$ are parameterized by prompts $\bm{\theta}=(\{\theta^{i}\}_{i=1}^{I},\theta^{a})$. By updating prompts via the risk-control component $M_{r}$, the whole system optimizes policies $\Pi_{\bm{\theta}}$ in a verbal reinforcement manner. By denoting daily profit & loss (PnL) by $R^{\Pi_{\bm{\theta}}}_{t}=\mathcal{R}(O_{t},B_{t},A_{t})$, the optimization objective for the whole system can be written as:
@@
which is a risk-sensitive optimization problem with respect to textual gradient descent, being essentially different from DRL algorithms for POMDP.
3.3 Textual Gradient-Descent
@@
4 General framework of FinCon
Figure 1: The general framework of FinCon.
In this section, we present the architecture of FinCon using a two-level hierarchy. First, we describe the hierarchical framework for coordinating the agents’ synchronous work and communication. Then, we elaborate on the functionalities of each module that constitutes each agent in FinCon.
4.1 Synthesized Multi-agent Hierarchical Structure Design
4.1.1 Manager-Analyst Agent Group
Analogous to human investment teamwork, FinCon establishes a unique hierarchy to organize the multi-agent system, synthesizing their efforts to achieve superior decision returns. The primary objective is to improve information presentation and comprehension while reducing unnecessary communication costs. The detailed working mechanism of each agent is explained and illustrated in Figure 2:
Analyst Agents. In FinCon, analyst agents distill concise investment insights from large volumes of multi-source market information for one specific trading target. To ensure high reasoning quality by reducing task load and enhancing analysis focus, each agent processes a single information source in a uni-modal manner, delivering pre-specified outputs based on prompts. This setup mirrors an effective human team structure, where each analyst specializes in a specific function and synchronously filters market noise to extract essential insights. These agents support the manager agent in making informed decisions by consolidating denoised investment information from multiple perspectives. We design seven distinct types of analyst agents using LLMs, each producing diverse investment insights, as shown in the upper part of Figure 2. Differentiated by input modalities, three textual data processing agents distill insights and sentiments from daily news and financial reports. An audio agent extracts investment tendencies from earnings call recordings using the Whisper API. Additionally, a data analysis agent and a stock selection agent compute key financial metrics, such as momentum and CVaR, based on tabular time series data. The stock selection agent also manages portfolio selection following the classic risk diversification method in quantitative finance [1], as detailed in Appendix. 7.5.
Manager Agent. In the FinCon system, the manager agent is the sole decision maker, generating trading actions for sequential financial tasks. For portfolio management, it computes portfolio weights using convex optimization techniques constrained by the direction of actions (see Equation. (3.1)). Four mechanisms support each decision: 1) Consolidating distilled investment insights from multiple analyst agents. 2) Receiving timely risk alerts and conceptual investment updates from the risk control component. 3) Refining investment beliefs about the impact of various information sources on the manager’s decisions regarding a specific trading target. 4) Retrieving self-reflections based on the reasoning outcomes of previous trading actions.
4.1.2 Risk-Control Component
We innovatively design a dual-level risk-control mechanism comprising within-episode and over-episode risk management. The within-episode risk control mechanism detects market risk within a single training episode, enabling the manager agent to adjust trading actions promptly to prevent potential losses, which takes account of short-term trading performance and market fluctuations. Additionally, the within-episode risk control mechanism also operates during the testing stage. In contrast, the over-episode risk control mechanism only functions during the training stage, and it provides updating directions for prompt optimization by comparing the trading performance of the current episode with the previous one. This reflection helps the manager agent update its investment beliefs based on the differences in performance. Informed by prior observations of market risk and profitability patterns, these two mechanisms can potentially avoid repetitive investment mistakes, thereby enhancing gains for future investments.
Within-Episode Risk Control: . The within-episode risk alert is triggered by an instant drop in CVaR value. The rationale of this setting is that Conditional Value at Risk (CVaR) typically represents the average of the bottom 1% of past daily trading Profits and Losses (PnLs). When the CVaR value decreases on a given day, it often means that the PnLs from our trading decisions on the previous day (or days) are in the bottom 1%, indicating a potentially high-risk market condition. Once the within-episode risk is initiated, the manager agent is instructed to adopt a risk-averse attitude for the current day’s trading actions, regardless of the prior risk status.
Over-Episode Risk Control: The over-episode investment belief updates facilitate timely adjustments in the emphasis placed on the analyst’s information distillation and the manager’s action generation. Through the Actor-Critic mechanism, FinCon episodically optimizes its investment strategy for a given trading target by deeply reflecting on continuous winning and losing actions. The episodic reflection of investment beliefs is generated by a unique Conceptual Verbal Reinforcement (CVRF). CVRF evaluates the performance of every two adjacent training episodes based on the information perspectives provided by analysts and available in the manager’s trading reasoning. It then conceptualizes and attributes the evaluation outcomes to these aspects. By comparing the conceptualized aspects of investment insights from both more profitable and less profitable episodes, the system informs the manager and analyst agents about the necessary belief adjustments for each market information aspect to prioritize for greater profit, following the steps of Algorithm 1. CVRF utilizes text-based gradient descent to provide optimal conceptual investment guidance on a given target asset, substantially refining prompts with the latest investment beliefs for the manager agent. The obtained guidance is also meticulously organized from perspectives corresponding to insights from specific analyst agents, key financial indicators such as historical momentum, or other necessary perspectives.
@@ 4.1.2 Risk-Control Component
These investment belief updates are initially received by the manager agent and selectively propagated to relevant agent nodes, thereby minimizing over-communication. Additionally, unlike the text-based gradient descent proposed by Tang et al.[19], which uses prompt editing distance as the learning rate, we obtain investment belief prompt updates by measuring the overlapping percentage of trading actions between two adjacent training trajectories achieved at each investment belief update. This approach has proven effective in enhancing the performance of a synthesized agent system, where each worker has a clearly defined and specialized responsibility.
@@
        Get the investment trajectory $\mathcal{H}_{k}$ and calculate objective function value (1).
@@
     Compare the objective function value 1 of episodes $k-1$ $k$, and decide which episode has higher performance;

### v3

@@ Abstract
Large language models (LLMs) have shown potential in complex financial tasks, but sequential financial decision-making remains challenging due to the volatile environment and the need for intelligent risk management. While LLM-based agent systems have achieved impressive returns, optimizing multi-source information synthesis and decision-making through timely experience refinement is underexplored. We introduce FinCon, an LLM-based multi-agent framework with conceptual verbal reinforcement for diverse financial tasks. Inspired by real-world investment firm structures, FinCon employs a manager-analyst hierarchy, enabling synchronized cross-functional agent collaboration towards unified goals via natural language interactions. Its dual-level risk-control component enhances decision-making by monitoring daily market risk and updating systematic investment beliefs through self-critique. These conceptualized beliefs provide verbal reinforcement for future decisions, selectively propagated to relevant agents, improving performance while reducing unnecessary peer-to-peer communication costs. FinCon generalizes well across tasks, including single stock trading and portfolio management.  [^1] We will release the code and demo in the following repo https://github.com/The-FinAI/FinCon
@@ 1 Introduction
The intricacies and fluctuations inherent in financial markets pose significant challenges for making high-quality, sequential investment decisions. In tasks such as single stock trading and portfolio management, each intelligent decision is driven by multiple market interactions and the integration of diverse information streams, characterized by varying levels of timeliness and modalities [1, 2]. The primary objective of these tasks is to maximize profit while managing present market risks in an open-ended environment.
@@
In practice, trading firms often depend on synthesized teamwork, structured hierarchically with functional roles such as data analysts, risk analysts, and portfolio managers communicating across levels [3, 4]. These roles are responsible for the careful integration of diverse resources. However, the cognitive limitations of human team members can hinder their capacity to rapidly process market signals and achieve optimal investment outcomes [5].
@@
To enhance investment returns and address the limitations of human decision-making, various studies have explored methods such as deep reinforcement learning (DRL) to develop agent systems that simulate market environments and automate investment strategies [6, 7, 8]. Concurrently, advancements in large language models (LLMs) have shown great potential in performing complex tasks, including reasoning [9, 10], tool-using [11], planning [12], decision-making [13, 14], and even in various financial applications [15, 16, 17, 18, 19], suggesting they may surpass existing agent architectures. Language agents, in particular, are distinguished by their human-like communication and flexible, prompt-based structures, making them well-suited to diverse decision-making settings [20, 21, 22, 23].
To achieve optimal decision-making performance, two critical factors must be considered: (1) Organizing agents to facilitate effective teamwork and efficient communication, and (2) Enabling agents to continuously learn and refine their actions. Studies have shown that mimicking human organizational structures can successfully coordinate language agents for specific tasks [24, 25, 26]. Additionally, recent advances in textual gradient-based prompt optimization [27, 28] and verbal reinforcement [29, 30] have proven effective in iteratively improving the reasoning and decision-making capabilities of language agents.
@@
Language agent systems designed for financial decision-making, such as FinGPT [31], FinMem [32], and FinAgent [33], have shown strong performance. However, they face several limitations. First, their reliance on agents’ risk preferences based on short-term market fluctuations fails to control long-term risk exposure, potentially overlooking fundamental factors driving investment returns. A more effective approach is to quantify investment risks using established measures of risk from quantitative finance [34, 35]. Second, these systems are often limited to single-asset trading tasks, making them less adaptable to multi-asset financial applications like portfolio management. Third, they place significant pressure on a single agent to understand and process information within a constrained context window, which can degrade decision quality. Although approaches like StockAgent [36] use multi-agent systems for stock trading, their reliance on extensive discussions between numerous LLM agents leads to high communication costs and slow decision-making. Moreover, the absence of a clear optimization objective can compromise outcome effectiveness. Additional related work in the literature is discussed in the Appendix A.1.
@@
To address these issues, we propose FinCon, an LLM-based multi-agent framework for critical financial tasks, such as single-stock trading and portfolio management, as shown in Figure 1. Our main contributions are: 1) Inspired by real-world investment roles, we introduce a novel Synthesized Manager-Analyst hierarchical communication structure with a risk-control component. This structure allocates financial data from different sources to corresponding functional analyst agents, allowing them to focus on specific insights, while the manager consolidates these inputs to make informed trading decisions. The streamlined communication reduces redundant peer-to-peer interaction, lowering costs and improving efficiency. 2) Our framework generalizes beyond stock trading to handle portfolio management, an area not previously addressed by other financial language agent systems. 3) We developed a dual-level risk control component to update risk assessments both within and across episodes. Within episodes, risk is supervised using the Conditional Value at Risk (CVaR), a quantile-based risk measure [37]. Across episodes, we introduced a verbal reinforcement mechanism, where investment beliefs are updated based on reasoning trajectories and profit-and-loss (PnL) trends, distilled into conceptual perspectives. These insights are selectively back-propagated from the manager to relevant analyst agents. Our ablation studies demonstrate the effectiveness of this risk control design in managing market risk and enhancing trading performance.
Figure 1: The general framework of FinCon.
@@ 2.1 Financial Decision-making Tasks Formulation
Single Stock Trading Tasks. FinCon uses analyst agents group $\{M_{pr}^{i}\}_{i=1}^{I}$ to process multi-modal market information sources. The processed information is then used by a manager agent $M_{a}$ to make trading decisions (buy, sell, hold), and to provide relevant reasoning texts. Note that the “sell” signal means the system makes a “short-selling” decision, that is, a negative trading position is allowed. Additionally, FinCon evaluates the daily investment risk, followed by prompt-optimization for the manager agent from risk-control component $M_{r}$.
Portfolio Trading Tasks. In addition to processing multi-modal market information, the analyst agents also construct a stock pool for portfolio management by considering the statistical correlations between stock returns. The manager agent then makes trading decisions for each stock in the pool. Finally, the manager agent determines the portfolio weights for all stocks using an external optimization solver that applies the mean-variance optimization described below [39]:
$\max_{\textbf{w}}\langle\textbf{w},\mu\rangle-\langle\textbf{w},\Sigma\textbf{w}\rangle\quad\text{s.t.}~{}w_{n}=\begin{cases}\in[0,1],&\text{``buy''}\\ \in[-1,0],&\text{``sell''}\\ =0,&\text{``hold''}\end{cases},~{}~{}\forall n\in\{1,\cdots,N\}$ (1)
where $\textbf{w}=(w_{1}\cdots,w_{N})\in\mathbb{R}^{N}$ is portfolio weights vector, $\mu$ and $\Sigma$ are the shrinkage estimators of $N$-dimensional sample expected return and $N\times N$ sample covariance matrix of chosen stocks’ daily return sequences respectively [35]. We note that portfolio weights are rebalanced on daily basis. In our implementation, we begin by calculating the portfolio weights through solving the aforementioned optimization problem. Next, the target positions are determined by linearly scaling these portfolio weights from the previous step.
@@ 2.2 Modeling Quantitative Trading as POMDP
Then, our multi-agent system is supposed to learn the policies of all agents: the policies of analyst agents $\pi_{\theta^{i}}^{i}:\mathcal{X}\to\mathcal{A}^{i},i\in\{1,\cdots,I\}$ (the ways to process information, i.e. $\hat{O}_{t}^{i}\sim\pi_{\theta^{i}}^{i}(\cdot|O_{t}^{i})$), and the policy of manager agent $\pi_{\theta^{a}}:\mathcal{A}\times\mathcal{Y}\to\mathbb{A}$ (the ways to make trading decisions, i.e. $A_{t}\sim\pi_{\theta^{a}}(\cdot|\hat{O}_{t},B_{t})$) such that the system maximizes cumulative trading reward while controlling risk [43]. All policies $\Pi_{\bm{\theta}}=(\{\pi_{\theta^{i}}^{i}\}_{i=1}^{I},\pi_{\theta^{a}})$ are parameterized by textual prompts $\bm{\theta}=(\{\theta^{i}\}_{i=1}^{I},\theta^{a})$. By updating prompts via the risk-control component $M_{r}$, the whole system optimizes policies $\Pi_{\bm{\theta}}$ in a verbal reinforcement manner. By denoting daily profit & loss (PnL) by $R^{\Pi_{\bm{\theta}}}_{t}=\mathcal{R}(O_{t},B_{t},A_{t})$, the optimization objective for the whole system can be written as:
@@
is a risk-sensitive optimization problem that leverages textual gradient descent, fundamentally differing from DRL algorithms designed for POMDPs. Further details on the textual gradient descent approach are provided in the Appendix A.2.
3 Architecture of FinCon
In this section, we present the architecture of FinCon using a two-level hierarchy. First, we describe the hierarchical framework for coordinating the agents’ synchronous work and communication. Then, we elaborate on the functionalities of each module that constitutes each agent in FinCon. Finally, we aim to elaborate on how FinCon solves the objective function expressed as Equation (2) through a verbal reinforcement approach.
3.1 Synthesized Multi-agent Hierarchical Structure Design
3.1.1 Manager-Analyst Agent Group
Analogous to human investment firm, FinCon establishes a unique hierarchical structure to organize its multi-agent system, synthesizing their efforts to achieve superior decision-making outcomes. The primary goal is to enhance information presentation and comprehension while minimizing unnecessary communication costs. The working mechanism of each agent is illustrated in Figure 2.
Analyst Agents. In FinCon, analyst agents distill concise investment insights from large volumes of multi-source market data, each focused on a specific trading target. To ensure high-quality reasoning by reducing task load and sharpening focus, each agent processes information from a single source in a uni-modal fashion, providing pre-specified outputs based on prompts. This setup mimics an efficient human team, where each analyst specializes in a specific function, filtering out market noise and extracting key insights. These agents assist the manager agent by consolidating denoised investment information from multiple perspectives. We implement seven distinct types of analyst agents using LLMs, each producing unique investment insights, as shown in the upper section of Figure 2. Based on input modalities, three textual data processing agents extract insights and sentiments from daily news and financial reports. An audio agent uses the Whisper API to interpret investment signals from earnings call recordings. Additionally, a data analysis agent and a stock selection agent compute critical financial metrics, such as momentum and CVaR, using tabular time series data. The stock selection agent also oversees portfolio selection by applying the classic risk diversification method in quantitative finance [1].
Manager Agent. In FinCon, the manager agent acts as the sole decision-maker, responsible for generating trading actions for sequential financial tasks. For portfolio management, it calculates portfolio weights using convex optimization techniques constrained by directional trading decisions (see optimization problem as presented in Formula (1)). Four key mechanisms support each decision: 1) Consolidating distilled insights from multiple analyst agents. 2) Receiving timely risk alerts and conceptual investment updates from the risk control component. 3) Refining its investment beliefs about the influence of different information sources on trading decisions for specific targets. 4) Conducting self-reflection by reviewing reasoning outcomes from previous trading actions.
3.1.2 Risk-Control Component
We have innovatively designed a dual-level risk-control mechanism consisting of within-episode and over-episode risk management. The within-episode mechanism detects market risk within a single training episode, allowing the manager agent to promptly adjust trading actions to mitigate potential losses by accounting for short-term trading performance and market fluctuations. This mechanism also operates during the testing phase. In contrast, the over-episode mechanism functions exclusively during the training stage, providing prompt optimization guidance by comparing the trading performance of the current episode with the previous one. This reflection enables the manager agent to update its investment beliefs based on performance differences. By drawing on prior observations of market risk and profitability patterns, these two mechanisms help avoid repeated investment errors, thereby enhancing future returns.
Within-Episode Risk Control: The within-episode risk alert is triggered by a sudden drop in the CVaR value. Conditional Value at Risk (CVaR) represents the average of the worst-performing 1% of daily trading Profits and Losses (PnLs). A decrease in CVaR typically indicates that recent trading decisions have led to PnLs within this bottom percentile, signaling a potentially high-risk market condition. When this occurs, the manager agent adopts a risk-averse stance for that day’s trading actions, regardless of the prior risk status.
Over-Episode Risk Control: The over-episode investment belief updates facilitate adjustments in the emphasis placed on analysts’ information distillation and the manager’s action generation. Through the Actor-Critic mechanism, FinCon episodically optimizes its investment strategy for a given trading target, as defined by objective (Equation (2)), by reflecting on a series of winning and losing actions. This episodic reflection is powered by a unique Conceptual Verbal Reinforcement (CVRF). CVRF assesses the performance of consecutive training episodes by analyzing the information perspectives provided by analysts and reflected in the manager’s decision-making. It then conceptualizes and attributes the evaluation outcomes to these specific aspects. By comparing the conceptualized insights from more profitable versus less profitable episodes, the system informs both the manager and analyst agents about necessary belief adjustments, helping prioritize the most relevant market information for increased profitability, as detailed in Algorithm 1. CVRF leverages text-based gradient descent to offer optimal conceptual investment guidance for the manager agent, refining prompts with the latest investment beliefs. The guidance is organized according to perspectives provided by the respective analyst agents, key financial indicators (such as historical momentum), or other crucial viewpoints.
@@ 3.1.2 Risk-Control Component
These belief updates are first received by the manager agent and then selectively propagated to relevant agents, minimizing over-communication. Unlike the text-based gradient descent proposed by Tang et al.[28], which uses prompt editing distance as a learning rate, we derive investment belief updates by measuring the overlapping percentage of trading actions between two consecutive training trajectories at each belief update, as presented in Table 1. This approach has proven effective in improving the performance of a synthesized agent system, where each worker has a clearly defined and specialized role. The above describes the workflow of FinCon during the training stage, while the workflow during the testing stage is detailed in the Appendix A.3.
@@
        Get the investment trajectory $\mathcal{H}_{k}$ and calculate the objective function value (Function (2)).
@@
     Compare the objective function values of episodes $k-1$ $k$, and decide which episode has higher performance;